A gopher has dug holes in opposite corners of a rectangular yard. If the yard is 7 meters by 7 meters, how far will the gopher have to run to get from one of its holes


1
Expert's answer
2021-10-20T18:28:47-0400

From the diagram above

We use the Pythagorean theorem

c= a2 + b2

where a=7, b=7 and c is unknown

c2 = 72 + 72

c2 = 49 + 49

c2 = 98

Finding the square root of both sides

"\sqrt{c²} = \sqrt{98}\\
c = 7\sqrt{2} m"